OpenStudy (anonymous):

Recall the equation for a circle with center (h,k) and radius r. At what point in the first quadrant does the line with equation y=2.5x+4 intersect the circle with radius 6 and center (0, 4)? =

OpenStudy (ranga):

Equation of a circle centered at (h,k) is: (x - h)^2 + (y - k)^2 = r^2 Put the values for h, k and r above. To find the point of intersection of the circle with the line y = 2.5x + 4, just set y = 2,5x + 4 in the circle equation and solve for x. You will get two values for x. Find the one that lies in the first quadrant. From the x value, put it in the straight line equation and find y.

OpenStudy (anonymous):

so is the circle equation (0-h)^2+(4-k)^2=6^2, if so where would I plug in the line equation?